A ball is thrown horizontally from the rooftop of a 150 m high building. If the ball

landed 17.5 m from the base of the building, what is the initial speed of the ball.


1
Expert's answer
2021-10-11T08:56:10-0400

Let's first find the time that the ball takes to reach the ground:


"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\cdot150\ m}{9.8\ \dfrac{m}{s^2}}}=5.53\ s."

Finally, we can find the initial speed of the ball:


"x=v_{0x}t,""v_{0x}=\dfrac{x}{t}=\dfrac{17.5\ m}{5.53\ s}=3.16\ \dfrac{m}{s}."